A 14.5 cm diameter loop of wire is initally oriented perpendicular to a 1.8 T magnetic field.The loop is rotated so that its plane is parallel to the field direction in 0.23 s.a.what is the average induced emf in the loop
Answer #1

Diameter of the loop d = 14.5 cm

Radius of the loop r = 7.25 cm = 7.25 *10^-2 m

Magnetic field B = 1.8 T

The loop rotated so that it's plane is parallel to the field direction in t = 0.23 sec

The average induced emf in the loop is

do A* B e dt

A = area of loop = π r^2 = π * ( 7.25*10^-2)^2

A = 0.0165 m^2

0.0165 1.8 e 0.23

e = emf = 0.129 V